If you are 16 years old now, which equation could be used to solve for your age, F, 15 years from now?
Answer by jim_thompson5910(35256)

Let x be your age now and F be your future age in 15 years.


(Age in 15 years) = (Age now) + 15

( F ) = ( x ) + 15

F = x + 15

F = 16 + 15 ... it says "If you are 16 years old now", so x = 16

F = 31

You'll be 31 years old 15 years from now.

The equation used is either F = x + 15 or F = 16 + 15 depending on how you do it. You can also manipulate the equation to get either F - x = 15 or F - 16 = 15. So there are many ways to do this.